Washington-Pacific (W-P) invests $3 million to buy a tract of land and plant some young pine trees. The trees can be harvested in 11 years, at which time W-P plans to sell the forest at an expected price of $6 million. What is W-P's expected rate of return? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

$6 mil = $3 mil * (1 + r)11

2 = (1 + r)11

1.0650 = 1 + r

r = 0.0650

r = 6.50%